Two resistors R 1 and R 2 are connected in parallel. Given R 1 = (5.0 ± 0.2) kΩ and R 2 = (10 ± 0.1) kΩ. The effective resistance is given by

Two resistors R1 and R2 are connected in parallel. Given R1 = (5.0 ± 0.2) kΩ and R2 = (10 ± 0.1) kΩ. The effective resistance is given by

Grade:11

1 Answers

Arun
6 years ago
First calculate the Nominal values of Resistances simply .i.i.e for 5 and 10 kilo ohm then calculate error separately
 
In parallel,
Rnet = R1*R2/(R1+R2) = 50/15 = 10/3 kilo ohm
 
 
for error take partial differnetiaion both sides.
d(Rnet) =d(R1) + d(R2) 
            = 0.2 + 0.1 = 0.3 kilo ohm
 
 
then net resistance with eroor is 10/3 + 0.3 k ohm
we always take maximum error thus with plus sign.